OpenStudy (brainzonly):

This is from one of my Brain Teasers books: My grandson is about as many days as my son in weeks, and my grandson is as many months as I am in years. My grandson, my son and I together are 120 years. Can you tell me my age in years ?

OpenStudy (brainzonly):

Since nobody will help me, I figured it out myself and I hope it's right! Let m be the grandmother's age in years. If s is the son's age in years, then the son is 52s weeks old. If g is the grandson's age in years, then the grandson is 365g days old. Thus, 365g = 52s. Since the grandson is 12g months old, 12g = m. Since the grandson, the son and the grandmother together are 120 years, g + s + m = 120. The above system of 3 equations in 3 unknowns (g, s and m) can be solved as follows. m / 12 + 365 m / (52 x 12) + m = 120 or 52 m + 365 m + 624 m = 624 x 120 or m = 624 x 120 / 1041 = 72. So, the grandmother is 72 years old.

OpenStudy (brainzonly):

@Kevin Solved another one by myself :(

OpenStudy (kevin):

XD